rxcklesskaisher:

A company makes triangular plates for individual slices of pizza. For each plate, the base is 5 inches and the height is 16 inches. The area of the top of the plate is ____inches squared.

rxcklesskaisher:

m answer s 80

rxcklesskaisher:

my is*

QuestionCoveBot:

@snowflake0531 ~ she can do this.

jhonyy9:

how you get this 80 ? what is formula for area of a triangle ?

rxcklesskaisher:

bxh

jhonyy9:

are of a triangle = b×h/2

rxcklesskaisher:

ok

rxcklesskaisher:

im lost

rxcklesskaisher:

so

rxcklesskaisher:

@jhonyy9

rxcklesskaisher:

is it 20

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher is it 20 \(\color{#0cbb34}{\text{End of Quote}}\) No, that is not correct  Area of a triangle = 1/2 *base * height the base is 5 inches the height is 16 inches area = 1/2 * 5 * 16 area = ??